A parallel‑plate capacitor has plates of 4.50 cm by 4.80 cm. The plates are separated by a 0.300-mm‑thick piece of paper. What is the capacitance C of this capacitor? The dielectric constant for paper is 2.1.
Answer #1

area of the plates=A=4.5*4.8=21.6 cm2=0.00216 m2

distance between the plates=d=0.3 mm=3*10-4 meters

k=dielectric constant=2.1

capacitance of a parallel plate capacitor is given as